Mind map
Lec 9 Continuous Probability

Map content
Lec 9 Continuous Probability
The Normal Distribution
- Probabilities for the Normal Probability Distribution
- Standardizing Normal Values = , by calculating its Z value:
- Using the Table of Cumulative Probabilities for a Standardized Value
- after finding the z score use the talbe in textbook to find the probability
- Using NORM.DIST
- NORM.DIST(value, mean, std deviation, True).
- The time to download a book to an e-reader is normally distributed, with a mean of 7 seconds, and a standard deviation of 2 seconds. What is the probability that you can download a book in 9 seconds or less? NORM.DIST(9,7,2,True)=.8413. What is the probability that the download will take more than 9 seconds? This will be 1- P (9 or less) so 1 - .8413 = .1587. What is the probability that the download will take more than 1 second? 1 – P(1 or less) = 1 – NORM.DIST(1,7,2,True)= .9987
- NORM.DIST(value, mean, std deviation, True).
- Probabilities for the Normal Probability Distribution
The Standard Normal Distribution
- Probabilities for the Standard Normal Distribution
- Using NORM.S.DIST = requires only the Z value.
- •Having found Z=1.00, it is possible to find the probability of observing a from NORM.S.DIST(1.00, True), which yields .8413.
- fine the z score
- use the text book table
- The time to download a book to an e-reader is normally distributed, with a mean of 7 seconds, and a standard deviation of 2 seconds. What is the probability that you can download a book in 9 seconds or less? Answer: First, find the Z value, and then consult the table. Answer: Find the Z Value: Z=(9-7)/2 = 1.00 Then consult the table: the probability is .8413.
- Using NORM.S.DIST = requires only the Z value.
- Probabilities for the Standard Normal Distribution
Finding an X Value, Given a Probability
- NORM.INV = requires (probability, mean, std dev
- : A certain distribution has a normal distribution, with mean = 7 and standard deviation=2. At what value of X is there .8413 probability of observations being less than X? NORM.INV(.8413,7,2) = 8.9996
- NORM.INV = requires (probability, mean, std dev
Finding a Z Value, Given a Probability
- NORM.S.INV()
Images and files in the map
Information
- Description
- Mind Map on Lec 9 Continuous Probability, created by ammar albaqshi on 17/02/2017.
- Created by
- ammar albaqshi
- Views
- 0
- Rating